We found that two of three PZT mirrors are at wrong place in the chambers.
Therefore we have to move these PZT mirrors together with their connections.
Here is a diagram for the current situation and the plan.
Basically mirror (A) must be associated to the output beam coming out from the SRM, but it was incorrectly put as a part of the input optics.
Similar to that, mirror (C) must belong to the input optics, but it is incorrectly being used as a part of OMC stuff.
Therefore we have to swap the positions of mirror (A) and mirror (C) as shown in the diagram above.
In addition to the mirror immigration, we also have to move their cables as well in order to keep the right functions.
We took a look at the length of the cables outside of the chambers in order to check if they are long enough or not.
And we found that the cables from c1asc (green line in the diagram) is not long enough, so we will put an extension D-sub cable.